OpenStudy (anonymous):

A box has 40 balls, of which 13 are red, 10 are blue, and 17 are yellow, as before. Balls are extracted at random but this time not returned to the box. What is the probability that after performing this test three times exactly two are red and one is yellow?

OpenStudy (anonymous):

use combinations

OpenStudy (anonymous):

@pgpilot326 u mean

OpenStudy (anonymous):



OpenStudy (anonymous):

i know that the fomular for combination is : \[\frac{ n! }{ r!(n-1) }\]

OpenStudy (anonymous):

for the denominator\(\left(\begin{matrix}40 \\ 3\end{matrix}\right)\) is appropriate what about the numerator?

OpenStudy (anonymous):